Calculate the gravitational force between two objects when they are 0.630m apart. Each object has a mass of 7.00kg. 


1
Expert's answer
2022-02-15T09:57:50-0500

Given:

"m_1=m_2=\rm 7.00\:\rm kg"

"r=0.630\:\rm m"


The law of universal gravitation says

"F=G\frac{m_1m_2}{r^2}=6.67*10^{-11}\frac{7.00*7.00}{0.630^2}=8.23*10^{-9}\:\rm N"